OpenStudy (anonymous):

Please help me with rate of change.. Your riding a bike. It takes you 36 minutes to go 11 miles. If you continue traveling at the same rate, how long will it take you to go 16 miles. (If necessary round to the nearest hundredth).

OpenStudy (texaschic101):

make a proportion.. 36/11 = x/16.....36 min to 11 miles = x min to 16 miles cross multiply (11)(x) = (36)(16) 11x = 576 x = 576/11 x = 52.363636...rounded to the nearest hundredth = 52.36